needhelpnowpls:

point M on a segment with endpoints X (1, −2) and Y (10, 3) partitions the segment in a 5:1 ratio. Find M. You must show all work to receive credit.

needhelpnowpls:

help pls

Chloeisdum:

@tetsxpreme

minustempo:

If a segment needs to be split into a 5:1 ratio, then essentially you split it into 6 equal parts, and so one side has 1 part, the other has 5 parts. So first, use the pythagorean theorem to find the distance between the two points.

needhelpnowpls:

how do i use the pythagorean theorem

minustempo:

a^2 + b^2 = c^2 a is the change in x, b is the change in y

needhelpnowpls:

sorry i just dont understand

mhanifa:

Since XM: MY = 5 : 1, to find the point M you need to add 5 parts out of 5 + 1 = 6 parts of the difference of the end points of the segment XY. Use the formula below to find coordinates of point M: x = 1 + 5/6(10 - 1) y = - 2 + 5/6(3 - (-2))

surjithayer:

let the coordinates of M be (x,y) if M divides A(x1,y1) and B(x2,y2) in the ratio m:n then \[x=\frac{ mx2+nx1 }{ m+n },y=\frac{ my2+ny1 }{ m+n}\] \[x=\frac{ 5\times 10+1\times 1}{ 5+1}=\frac{ 50+1 }{ 6 }=\frac{ 51 }{ 6 }=\frac{ 17 }{ 2 }\] \[y=\frac{ 5\times3+1\times -2 }{ 5+1}=\frac{ 15-2 }{ 6 }=\frac{ 13 }{ 6 }\] so coordinates of M are \[(\frac{ 17 }{ 2 },\frac{ 13 }{ 6})\]